# Exploring the 6G Potentials: Immersive, Hyper Reliable, and Low-Latency Communication

### v1

@@ III-A2 RIS for LoS connectivity and diversity
The RIS technology presents a groundbreaking advancement in wireless communication, offering an evolutionary approach to signal propagation and control. RISs are artificial surfaces with numerous small, electronically controlled passive elements capable of altering the phase and amplitude of the incoming signal, thus enabling the redirection or reshaping of the signal path[9]. Therefore, an RIS enables a favorable and adaptable wireless environment for signal propagation, which can enhance the performance of wireless networks from several perspectives. First, considering the ever-increasing demand for ubiquitous seamless connectivity, wireless networks require deploying more base stations to ensure reliable connectivity for users. However, this will bring a higher energy consumption, higher deployment and maintenance costs, and a higher level of interference to deal with. RIS passive elements and their light weight make them capable of being deployed in many strategic spots in the city at a relatively low cost, which enables them to overcome the communication distance limitations of the THz frequency band by offering LoS links, thus paving the way for IHRLLC.
Second, RISs can be strategically positioned to create a scattering channel, which is desirable for enhancing the performance of umMIMO systems in THz communications. As a result, RIS systems can add a degree of freedom to umMIMO systems and improve the reliability of wireless networks.
@@
III-B Non-terrestrial networks
@@
Non-terrestrial networks (NTN), which include space satellites and aerial platforms such as high altitude platform stations (HAPS) and unmanned aerial vehicles (UAVs), present a promising architecture for 6G wireless networks. Traditionally, NTN have been considered a potential technology for enhancing wireless networks coverage in rural and remote areas. However, given the ever-increasing capacity demands of future wireless networks and the need for a resilient and reliable network architecture, the new paradigm of employing the NTN is to integrate them into urban terrestrial wireless networks to guarantee an immersive and hyper reliable connection for the users[12]. HAPS, located at an altitude of about $20$ Km above the ground, can provide a one-way propagation delay of the order of microseconds, which is much less than that of the satellite networks. In addition, as depicted in Fig. 3, HAPS can play the role of backhauling for the UAVs via enabling airborne integrated access and backhauling (IAB). In applications such as UAM, multi-connectivity through aerial platforms, such as UAVs and HAPS, will be a potential solution in case of losing connection with either the main base station or the RIS (Fig. 3).
@@
In the previous sections, various enabling technologies were discussed to facilitate IHRLLC. The implementation of such sophisticated technologies, as well as the proper design of the corresponding parameters, is a critical challenge. Considering the umMIMO antennas and RISs, with a large number of elements along with the challenges of THz communication, designing the optimum parameters through traditional approaches incurs a high computational cost. To this end, AI algorithms can potentially produce revolutionary changes in future wireless networks and overcome computational challenges. In the following discussion, we investigate promising novel computational-related approaches, including learn-to-optimize (L2O), generative AI (GenAI), and quantum computing technologies as methods to solve the aforementioned computational challenges in 6G wireless networks.
III-C AI technologies: L2O and GenAI
@@
L2O leverages machine learning (ML) to improve the current optimization algorithms or develop new optimization approaches [10]. Conventional optimization techniques often rely on fundamental methods such as gradient descent, conjugate gradient, and Newton steps, which are theoretically robust and straightforward. These methods are chosen based on the type of problem and are expected to provide solutions that meet their theoretical guarantees. However, L2O presents a different approach in which optimization methods are developed through training based on their performance with sample problems (Fig. 4a). These methods may not have a solid theoretical foundation, but are designed to enhance performance through training, which usually occurs online. At the same time, this training aims to make the actual application of the methods more efficient. In situations involving complex problems such as nonconvex optimization or inverse problems, a well-trained L2O method might deliver better results than traditional methods.
@@
Although conventional AI/ML algorithms have already been utilized in developing algorithms for wireless networks, GenAI has recently been considered a potential tool for future wireless networks. GenAI can be employed in wireless networks to overcome the challenges of the conventional AI algorithms[11]. Specifically, GenAI models can be utilized in various physical layer aspects of the networks, such as IRS deployment design, joint source-channel coding, beamforming, etc., as illustrated in Fig. 4b. In addition, the GenAI models can generate data for training purposes and can be implemented near the training locations.
@@ III-D Quantum Computing
Future wireless networks will experience massive communications among users in various usage scenarios. This makes it increasingly difficult to decide the order in which resource allocation should be made. These challenges can be defined as combinatorial problems which involve determining the service order for different users and the routes for traffic transfer to ensure everyone receives the promised QoE. Additionally, networks are becoming software-oriented through the use of the SDN and NFV technologies. Next-generation networks will employ AI and cognitive technologies for self-healing, self-management, and decision making, which require massive computing capabilities, leading to high latency [13]. Traditionally, combinatorial problems have so far been largely solved using heuristic methods that simplify the problem. However, these methods typically yield approximate answers that, in certain situations, may lead to inferior performance. Optimization methods based on statistics and data, on the other hand, consume a significant amount of energy and may not be suitable for dynamic situations that require quick reactions. To address these issues, hybrid algorithms combining quantum and classical computing, such as the quantum approximate optimization algorithm (QAOA), are used. These algorithms can be deployed at the core network, in various scales according to the hardware capabilities, offering considerable performance improvements in many combinatorial optimization problems [14]. Hybrid algorithms optimize the strengths of both computing approaches, ultimately increasing the network capacity. This ensures that demanding applications receive the necessary service quality, even in a busy network, while also saving energy.
Figure 5: NDT-empowered network for IHRLLC in mobile tele-surgery and UAM use cases.
@@ III-E Network Digital Twin (NDT)
Digital twin (DT) is an emerging technology that virtually represents a physical system. It accurately models complex systems by replicating real physical entities and functionalities of a physical system into the digital world in real-time. Combining the DT technology with AI/ML tools, real-time data can be used to improve the operation and performance of the physical system in terms of precise decision-making and efficient resource utilization. The network DT (NDT) has recently been proposed as a key enabler for efficient control and management of modern communication networks [15]. Through the NDT dynamic state of the network, the status of the edge and cloud servers can be captured in real-time. Furthermore, combining 3D map services with NDT can determine the location and health of network infrastructure elements such as base stations and RISs. The NDT does therefore improve network performance. However, it also suffers from problems of data asynchronization and deviation between the NDT and the real network. Such issues can be addressed by appropriately applying the tools from the AI/ML framework.
@@ IV Proposed Network Architecture for IHRLLC
We now propose a network architecture that integrates the aforementioned technologies to support and achieve the requirement for the IHRLLC service class. We use NDT as the underlying technology that helps the IHRLLC use cases to make more informed and swift decisions. In particular, NDT is used to gain information about the infrastructure locations and the status of edge and cloud computing servers in the entire network. As an example, the proposed NDT-empowered network for IHRLLC in mobile tele-surgery and UAM use cases is shown in Fig. 5. Essentially, the network relies on the THz band to support immersive communication. To reliably transmit data at high data rates, the communication relies on numerous RISs deployed at different locations throughout the city. After obtaining the relative location information via NDT, and with the help of the umMIMO, the ambulance continuously forms pencil beams toward multiple RISs in LoS with an ambulance. The use or multiple RISs not only facilitates LoS communications but also enhances diversity reception. Both the LoS link and diversity reception significantly improve reliability. Each RIS is tuned such that all reflected signals are directed to the nearest base station and subsequently to the core network and then eventually to the hospital. Furthermore, multi-connectivity transmission is available. If the ambulance can not find an RIS in the range, it then starts communication via a dedicated UAV that is associated with this ambulance. For the UAM use case, UAVs and HAPS will be part of the IHRLLC network, which ensures seamless connectivity for the user.
AI enablers (L2O and GenAI) and quantum computing play a crucial role in the proposed network by reducing computational delays for algorithms related to efficient resource allocation, beamforming design, beam-tracking, AL/ML-based operations, and various functions for implementing SDN and NFV. The L2O and GenAI models can be utilized at the base stations and the core network, in different scales, to support designing efficient algorithms. Accordingly, quantum computing resources are distributed across the core network components. In this way, tasks requiring heavy computation are delegated to the core network due to the higher processing capabilities. On the other hand, tasks requiring real-time decision making will be handled at the base stations using L2O and GenAI models. Task allocation is based on the status of all the servers, monitored via NDT. Additionally, HAPS serves as a computing platform for specific use cases such as UAM.
@@ V Challenges and Open Issues
In the previous sections, we discussed the 6G use cases that require IHRLLC to achieve an acceptable QoE, and we proposed a combination of technologies to achieve IHRLLC. However, designing an E2E wireless network, equipped with the proposed technologies, and is capable of dealing with a large volume of data in a hyper reliable manner within a hyper low-latency, is a significant challenge. Considering the THz frequency band, combined with umMIMO and RIS technologies, the development of hardware devices such as antenna elements and electronic or photonic devices that can generate high frequency electromagnetic waves is an open area of research among wireless researchers. The THz-specific air interface techniques and multiple access protocols need further exploration to ensure network reliability. Furthermore, the human health and safety issues associated with the THz technology should be thoroughly analyzed and evaluated.
Although AI has been introduced as a game changer for future wireless networks, specific characteristics of these networks make the implementation of the AI algorithms challenging. One of the key challenges is the randomness of the wireless network environments inherited from the random channel models. This is considerably important in urban areas where wireless channels experience significant small-scale fading. Consequently, developing AI-driven algorithms to support the E2E network functionality in a robust, resilient, dynamic, and efficient manner is a significant challenge and, hence, an important research area. This is in addition to the privacy and standardization concerns related to the implementation of AI algorithms. Finally, integrating a combination of all potential technologies into wireless networks necessitates interdisciplinary research among experts in wireless networks, AI, and quantum computing to ensure the efficiency and scalability of the developed algorithms.
@@ VI Conclusion
In this paper, we discussed the 6G use cases that require achieving immersive, hyper reliable, low-latency communication (IHRLLC) simultaneously. Accordingly, we proposed a new class, termed the IHRLLC class, to be considered as one of the main 6G service classes. We also proposed potential network solutions to meet the KPI requirements and overcome the challenges of achieving multiple KPIs. To this end, THz communication, combined with umMIMO and RIS, plays an important role in making IHRLLC feasible. An NDT-empowered network architecture was proposed to facilitate the IHRLLC applications. Moreover, novel computation and communication techniques, including L2O, GenAI, quantum computing, and NDT, were proposed to design the parameters of various parts of the architecture.
@@
Afsoon Alidadi Shamsabadi [M] (afsoonalidadishamsa@sce.carleton.ca) is a PhD Candidate in the Non-Terrestrial Networks (NTN) Lab in the Department of SCE at Carleton University, Canada.
Animesh Yadav [SM] (yadava@ohio.edu) is an assistant professor in the School of EECS at Ohio University, Athens, OH, USA.
Yasser Gadallah [SM] (ygadallah@aucegypt.edu) is a Professor at the Department of ECE at The American University in Cairo (AUC), Egypt. His research interests include IoT, wireless communication and smart systems. He has extensive telecommunications industrial experience, having worked at several high-profile industry-leading companies in this area.
Halim Yanikomeroglu [F] (halim@sce.carleton.ca) is a Chancellor’s Professor of Systems and Computer Engineering in the Non-Terrestrial Networks Lab., at Carleton University, Canada. He is a Fellow of EIC (Engineering Institute of Canada), CAE (Canadian Academy of Engineering), and AAIA (Asia-Pacific Artificial Intelligence Association); and a Distinguished Speaker for IEEE Communications Society and IEEE Vehicular Technology Society.

### v2

@@ III-A2 RIS for LoS connectivity and diversity
The RIS technology presents a groundbreaking advancement in wireless communication, offering an evolutionary approach to signal propagation and control. RISs are artificial surfaces with numerous small, electronically controlled passive elements capable of altering the phase and amplitude of the incoming signal, thus enabling the redirection or reshaping of the signal path [9]. Therefore, an RIS enables a favorable and adaptable wireless environment for signal propagation, which can enhance the performance of wireless networks, specifically those operating within the THz frequency band. RISs can be strategically positioned to create a scattering channel, which is desirable for enhancing the performance of umMIMO systems in THz communications. As a result, RIS systems can add a degree of freedom to umMIMO systems and improve the reliability of wireless networks. In addition, considering the IHRLLC service class use cases that require a seamless guaranteed level of KPIs, wireless networks require deploying more base stations to ensure reliable connectivity for users. However, this will bring higher energy consumption, higher deployment and maintenance costs, and a higher level of interference to deal with. RIS passive elements and their light weight make them capable of being deployed in many strategic spots in the city at a relatively low cost, which enables them to overcome the communication distance limitations of the THz frequency band by offering LoS links, thus paving the way for IHRLLC.
@@
III-B Multi-Connectivity through Non-Terrestrial Networks
@@
Non-terrestrial networks (NTN), which include space satellites and aerial platforms, such as high altitude platform stations (HAPS) and unmanned aerial vehicles (UAVs), present a promising architecture for 6G wireless networks. Traditionally, NTN have been considered a potential technology for coverage enhancement in rural and remote areas. However, given the ever-increasing capacity demands of future wireless networks and the need for a resilient and reliable network architecture, the new paradigm of employing the NTN is to integrate them into urban terrestrial wireless networks to guarantee an immersive and hyper reliable connection for the users [12]. To this end, aerial platforms have the potential to be part of the proposed architecture for IHRLLC service class. As depicted in Fig. 3, in applications, such as UAM, multi-connectivity through aerial platforms, such as UAVs and HAPS, will be a potential solution in case of losing connection with either the main base station or the RIS. Moreover, HAPS, located at an altitude of about $20$ Km above the ground, can provide a one-way propagation delay of the order of microseconds, which is much less than that of the satellite networks. In addition, as depicted in Fig. 3, HAPS can play the role of backhauling for the UAVs by enabling airborne integrated access and backhauling (IAB).
@@
In the previous sections, various enabling technologies were discussed to facilitate IHRLLC. The implementation of such sophisticated technologies, as well as the proper design of the corresponding parameters, is a critical challenge. Considering the umMIMO antennas and RISs, with a large number of antenna elements along with the challenges of THz communication, designing the optimum parameters through traditional approaches incurs a high computational cost. To this end, AI algorithms can potentially produce revolutionary changes in future wireless networks and overcome computational challenges. In particular, AI can be integrated with wireless networks as a distinct network function or be embedded within existing functions, depending on network demands. In scenarios where AI needs to handle complex, centralized tasks, AI can be deployed as a dedicated network function, responsible for high-level decision-making, and coordination with other functions to optimize the overall network performance. Alternatively, AI can be integrated as a module within existing network functions to enhance their operations. This modular integration enhances the intelligence and adaptability of current architectures without introducing a completely new function. In the following discussion, we investigate promising novel computational-related approaches, including learn-to-optimize (L2O), generative AI (GenAI), and quantum computing technologies as methods to solve the aforementioned computational challenges in 6G wireless networks.
III-C AI Enablers: L2O and GenAI
@@
L2O leverages machine learning (ML) to improve current optimization algorithms or develop new optimization approaches [10]. Conventional optimization techniques often rely on fundamental methods, such as gradient descent, conjugate gradient, and Newton steps, which are theoretically robust and straightforward. These methods are chosen based on the type of problem and are expected to provide solutions that meet their theoretical guarantees. However, L2O presents a different approach in which optimization methods are developed through training based on their performance with sample problems (Fig. 4a). These methods may not have a solid theoretical foundation but are designed to enhance performance through training, which usually occurs online. At the same time, this training aims to make the actual application of the methods more efficient. In situations involving complex problems, such as nonconvex optimization or inverse problems, a well-trained L2O method might deliver better results than traditional methods.
@@
Although conventional AI/ML algorithms have already been utilized in the development of algorithms for wireless networks, GenAI has recently been considered a potential tool for future wireless networks. GenAI can be employed in wireless networks to overcome the challenges of the conventional AI algorithms [11]. Specifically, GenAI models can be utilized in various physical layer aspects of the networks, such as RIS deployment design, joint source-channel coding, beamforming, etc., as illustrated in Fig. 4b. In addition, GenAI models can generate data for training purposes and can be implemented near training locations.
@@ III-D Quantum Computing
Next-generation wireless networks will experience massive communications among users in various usage scenarios. This significantly complicates the decision-making process for determining the order of resource allocation. Additionally, these networks will leverage AI and cognitive technologies for self-healing, self-management, and decision making, requiring massive computing capabilities [13]. These challenges can be framed as combinatorial problems which involve determining the service order for different users and the routes for traffic transfer to ensure that everyone receives the promised QoE.
Traditionally, combinatorial problems have been largely solved using heuristic methods that simplify the problem. However, these methods typically yield approximate answers that, in certain situations, may lead to inferior performance. On the other hand, optimization methods based on statistics and data, while more accurate, are energy-intensive and unsuitable for dynamic situations requiring quick responses. To address these limitations, hybrid algorithms, like the quantum approximate optimization algorithm (QAOA), that combine quantum and classical computing, offer a promising solution. These algorithms can be deployed in the core network at various scales, according to hardware capabilities, offering considerable performance improvements in solving combinatorial optimization problems [14]. Hybrid algorithms ensure that demanding applications receive the necessary service quality, even during peak network usage, and also contribute to energy saving.
@@ III-E Network Digital Twin (NDT)
Digital twin (DT) is an emerging technology that virtually represents a physical system. It accurately models complex systems by replicating real physical entities and functionalities of a physical system into the digital world in real-time. Combining the DT technology with AI/ML tools, real-time data can be leveraged to improve the operation and performance of the physical system through precise decision-making and efficient resource utilization. Recently, the network DT (NDT) has been proposed as a key enabler for efficient control and management of modern communication networks [15]. NDT enables real-time monitoring of the status of edge and cloud servers. Additionally, integrating 3D map services with NDT allows for the determination of the location and health of network infrastructure elements, such as base stations and RISs. In this way, NDT significantly enhances network performance. However, it faces challenges, such as data asynchronization between the NDT and the real network. These challenges can be mitigated through the strategic application of AI/ML tools.
Figure 5: NDT-empowered network architecture for IHRLLC service class in 6G.
@@ IV Proposed Network Architecture for IHRLLC
We now propose a novel network architecture that integrates the aforementioned technologies to meet the requirements for the IHRLLC service class. The proposed NDT-empowered architecture, designed for the mobile tele-surgery and the UAM use cases, is shown in Fig. 5. This architecture benefits end users by enabling IHRLLC service while offering mobile operators and industries opportunities to enhance network architecture and develop new business models with improved QoE. In this architecture, the base stations are equipped with umMIMO antenna and communicate using the THz band to enable immersive communications. Given the short range communication nature of THz frequency band, numerous RISs are deployed at different locations throughout the city to reliably transmit data at high data rates. We use NDT as the underlying technology that helps the IHRLLC use cases to make more informed and swift decisions. Obtaining the relative location information via NDT, and with the help of the umMIMO, the user continuously forms pencil beams toward multiple RISs in LoS with the user. The use of multiple RISs not only facilitates LoS communications, but also enhances diversity reception, which significantly improves reliability. Accordingly, each RIS is pre-tuned so that all reflected signals are directed to the nearest base station and subsequently to the core network. Furthermore, the proposed architecture provides multi-connectivity transmission through aerial platforms for the users. For instance, in the mobile tele-surgery use case, if the ambulance can not find an RIS in the range, it then establishes a connection through a dedicated UAV that is associated with this ambulance. For the UAM use case, UAVs and HAPS will be part of the proposed IHRLLC architecture, ensuring seamless connectivity for the user.
AI enablers (L2O and GenAI) and quantum computing play a crucial role in the proposed network architecture by reducing computational delays for efficient resource allocation, beamforming design, beam-tracking, and various management functions. The L2O and GenAI models can be utilized at base stations, edge network (near the base station), and the core network, on different scales, to support the design of efficient algorithms. Accordingly, quantum computing resources are distributed across the edge nodes and core network components. In this way, tasks requiring heavy computation are delegated to the core network due to the higher processing capabilities. On the other hand, tasks requiring real-time decision making will be handled at the base stations using L2O and GenAI models. Task allocation is based on the status of all the servers, monitored via NDT. Additionally, in the proposed architecture, HAPS can serve as a powerful computing platform to offload the heavy non-real time computations from the base stations.
@@ V Challenges and Open Issues
Previous sections discuss the 6G use cases that require IHRLLC, and propose a combination of technologies and a unified network architecture to achieve IHRLLC with an acceptable QoE. However, designing an E2E wireless network capable of catering to IHRLLC use cases poses a significant challenge. Particularly, the developed algorithms in the proposed network architecture, namely, the resource allocation, beamforming design, user association, etc., should be capable of dealing with large-scale network parameters, related to the ultra massive number of antenna elements and RIS components. As a result, the scalability of the developed AI-driven algorithms is a critical challenge. On the other hand, although the IHRLLC service class deals with the use cases that do not involve a large number of users, the algorithms should be scalable from the number of users perspective in order to ensure the expected QoE. In addition, efficient interference management approaches should be employed in the network to manage and mitigate the propagated interference among the users. In addition to the challenges related to scalability, efficient implementation of the THz frequency band, combined with umMIMO and RIS technologies, necessitates the development of hardware devices (antenna elements and electronic/photonic devices) that can generate high frequency electromagnetic waves. This challenge is an open area of research among wireless researchers. In addition, the THz-specific air interface techniques and multiple access protocols need further exploration to ensure network reliability. Furthermore, human health and safety issues, associated with the THz technology, should be thoroughly analyzed and evaluated. To address this issue, regulatory policies should be considered while implementing the THz frequency band.
Although AI has been introduced as a game changer for future wireless networks, the specific characteristics of the wireless networks make the implementation of AI algorithms challenging. One of the key challenges is the randomness of the wireless network environments, inherited from the random channel models. This is considerably important in urban areas where wireless channels experience significant random small-scale fading. This is in addition to the privacy and standardization concerns related to the implementation of AI algorithms. Finally, integrating a combination of all potential technologies into wireless networks necessitates interdisciplinary research among experts in wireless networks, AI, and quantum computing to ensure the efficiency and scalability of the developed algorithms. With a proper interdisciplinary coordination, the proposed IHRLLC service class and NDT-empowered network architecture have the potential to be part of the 6G standardization.
@@ VI Conclusion
In this paper, we proposed a new service class for 6G, referred to as immersive, hyper reliable, and low-latency communication (IHRLLC). The 6G use cases under the IHRLLC service class require achieving immersive throughput, with an envisioned peak data rate of $1$ Tbps, hyper-reliability, in the range of $10^{-5}$ to $10^{-7}$, and low-latency, with values between $0.1$ and $1$ ms, simultaneously. To achieve this, we proposed an NDT-empowered network architecture which incorporates potential technologies to meet the KPI requirements and overcome the challenges of simultaneously achieving multiple KPIs. To this end, THz communications, along with umMIMO and RISs, play an important role in enabling the IHRLLC class. In addition, novel computation and communication techniques, including L2O, GenAI, and quantum computing, were proposed to facilitate the design of the parameters of various parts of the architecture.
@@
Afsoon Alidadi Shamsabadi [SM] (afsoonalidadishamsa@sce.carleton.ca) is a PhD Candidate in the Non-Terrestrial Networks (NTN) Lab in the Systems and Computer Engineering Department at Carleton University, Canada. Her research focuses on the innovative integration of High Altitude Platform Stations (HAPS) with terrestrial wireless networks.
Animesh Yadav [SM] (yadava@ohio.edu) is an assistant professor in the School of EECS at Ohio University, Athens, OH, USA. He is the Senior Editor for IEEE Communications Letters and Associate Editor for Frontiers in Communications and Networks.
Yasser Gadallah [SM] (ygadallah@aucegypt.edu) is a Professor at the Department of ECE at The American University in Cairo (AUC), Egypt. His research interests include IoT, wireless communication, and smart systems. He has extensive telecommunications industrial experience, having worked at several high-profile industry-leading companies in this area.
Halim Yanikomeroglu [F] (halim@sce.carleton.ca) is a Chancellor’s Professor of Systems and Computer Engineering in the Non-Terrestrial Networks Lab, at Carleton University, Canada. His group’s focus is the wireless access architecture for the 2030s and 2040s, and non-terrestrial networks. He is a Fellow of EIC (Engineering Institute of Canada), CAE (Canadian Academy of Engineering), and AAIA (Asia-Pacific Artificial Intelligence Association); and a Distinguished Speaker for IEEE ComSoc and IEEE VTS.